OpenStudy (anonymous):

Clear fractions or decimals, solve and check. 0.91-0.2z = 1.23-0.6z

OpenStudy (cwrw238):

add 0.2z to both sides of equation then subtract 1.23 from both sides finally, divide both sides by coefficient of z

jhonyy9 (jhonyy9):

0,6z-0,2z=1,23-0,91 0,4z=2,14 z=2,14/0,4 =214/40 =107/20 =10,7/2 =5,35

OpenStudy (anonymous):

check maybe can help you http://www.wolframalpha.com/

OpenStudy (anonymous):

6x